Consider forest soil in contact with water containing 3.9 μg/L. Use the Freundlich relation to determine...

  1. Consider forest soil in contact with water containing 3.9 μg/L. Use the Freundlich relation to determine the equilibrium concentration of adsorbed cadmium. Assume KF = 0.0324 L/g and nF= 0.82. Why is it appropriate to use the Freundlich relation in this case?
Answer #1

7 m = 8.999/2 = 3,9x106g/ kg = 0.0324 4/gm 0.82 ng= 0.82 c=? clag r = legk th lefe cheg 3,98106 = lug 0.0324 + 2 loge lg 3.94